### v1

# Survey of Graph Neural Network for Internet of Things and NextG Networks

Abstract
The exponential increase in Internet of Things (IoT) devices coupled with 6G pushing towards higher data rates and connected devices has sparked a surge in data. Consequently, harnessing the full potential of data-driven machine learning has become one of the important thrusts. In addition to the advancement in wireless technology, it is important to efficiently use the resources available and meet the users’ requirements. Graph Neural Networks (GNNs) have emerged as a promising paradigm for effectively modeling and extracting insights which inherently exhibit complex network structures due to its high performance and accuracy, scalability, adaptability, and resource efficiency. There is a lack of a comprehensive survey that focuses on the applications and advances GNN has made in the context of IoT and Next Generation (NextG) networks. To bridge that gap, this survey starts by providing a detailed description of GNN’s terminologies, architecture, and the different types of GNNs. Then we provide a comprehensive survey of the advancements in applying GNNs for IoT from the perspective of data fusion and intrusion detection. Thereafter, we survey the impact GNN has made in improving spectrum awareness. Next, we provide a detailed account of how GNN has been leveraged for networking and tactical systems. Through this survey, we aim to provide a comprehensive resource for researchers to learn more about GNN in the context of wireless networks, and understand its state-of-the-art use cases while contrasting to other machine learning approaches. Finally, we also discussed the challenges and wide range of future research directions to further motivate the use of GNN for IoT and NextG Networks.
Index Terms:
Internet of Things, Next-Generation Networks, 6G, Graph Neural Networks, Survey, Applications
I Introduction
Internet of Things (IoT), generally defined as a network of interconnected devices, sensors, and other resources that communicate with each other over the internet. A simple IoT network consists of three layers namely the Perception or Sensing Layer, the Network Layer, and the Application Layer. The role of Perception or Sensing Layer is to gather the data from connected sensors or actuators such as air quality monitoring sensors, motion sensors, cameras, and temperature sensors, to name a few. The Network Layer is responsible for transmitting the collected data to the cloud or server through a centralized gateway. Examples of network layer protocols include ZigBee, Bluetooth, Long Range (LoRa), Wireless Fidelity (WiFi), Near-Field Communication (NFC), etc. The cloud (or server) takes care of storing and transforming sensor raw data into human-understandable values as well as performing statistical analysis. The Application Layer provides access to the processed sensor data as well as enables interaction between the user and the sensors. In a more sophisticated IoT network, there is another layer between the network and application layer called the Processing Layer.
Figure 1: Applications of IoT.
I-A Applications of IoT
Over the past decade, the number of connected IoT devices has grown significantly. There are over 30 billion IoT devices, which are expected to quadruple by 2030 [1]. This can be attributed to its myriad of features such as improved quality of life [2], enhanced decision making [3], safety and security [4], efficiency and automation [5], improved accessibility [6], business innovation and growth [7], environmental sustainability [8]. These features have opened the door to the application of IoT across various domains in everyday life as shown in Figure 1. For example, IoT finds its application in smart home systems [9, 10], smart city [11, 12], smart energy systems [13, 14], smart infrastructure [15, 16, 17], smart logistics [18, 19], smart laboratory [20, 21], smart health [22, 23], smart factory [24, 25], smart campus [26], smart surveillance [27, 28], smart agriculture [29, 30], smart transportation [31, 32].
Figure 2: Vision of GNN-enabled 6G Wireless Network.
I-B Importance of 6G and NextG Networks
While IoT enables a wide range of applications for consumers, it has a significant effect on the underlying network and spectrum (i.e., IoT network layer) including but not limited to network and resource management, adversarial attacks, security and privacy concerns, and reliable quality of service (QoS). While 5G technology provided much-needed improvement over previous generations, the number of devices requiring enhanced capacity, ultra-low latency, and very high throughput continues to increase. To overcome these challenges, the research community has shifted focus to the 6G network as a necessary upgrade to match the user requirements [33, 34, 35, 36]. Compared to its predecessor, the 6G network is expected to have 10 times the peak data rate and 1 ms latency [33]. It is also expected that the 6G system will have 1000 times higher simultaneous wireless connectivity compared to the 5G system [33]. Millimeter wave (mmWave) and Terahertz (THz) bands will play a key role in enabling the high data rate and ultra-low latency communication links [37, 38, 39, 40, 41]. Compared to traditional cellular and WiFi networks that use 300 MHz to less than 30 GHz, mmwave will operate at 30–300 GHz while THz will operate at 0.1 THz to 10 THz. While we move from 5G to 6G, it is important to effectively re-use the existing infrastructure instead of the need to re-deploy new hardware and other resources. Open Radio Access Network (O-RAN) is considered the solution to create more open, flexible, and interoperable mobile networks [42]. O-RAN introduces a set of open, standardized interfaces to interact, control, and collect data from every node in the network. O-RAN consists of RAN Intelligent Controllers (RICs) that execute third-party applications to control the RAN functionalities [43]. Common RAN functionalities include resource management, mobility management, and QoS management to name a few.
I-C Era of Machine Learning
In the recent decade due to the exponential increase in computational resources and availability of digital data, Machine Learning (ML) has been making a significant impact in the research community [44, 45, 46]. This is due to the vast potential of ML techniques such as automation of tasks [47], enhanced user experience [48], resource-efficient big data transfer [49] and intelligent decision making [50]. ML is also expected to enable improved energy efficiency and reduced human intervention [51], secure 6G networks [52], intelligent 6G networks [53]. To this end, researchers have used different machine learning techniques such as Feed forward networks [54], Convolutional Neural Networks (CNN) [55, 56, 57], Gated Recurrent Units (GRU) [58, 59], Echo State Networks (ESN) [60, 61], Long-Short Term memory (LSTM) [62, 63] and even hybrid architectures [64]. Additionally, there are techniques such as Generative Adversarial Networks [65], Autoencoder [66], and a category by itself referred to as Deep Reinforcement Learning [67, 68, 69, 70, 71] that using a machine of these machine techniques as its function approximator. These ML techniques fall under a general category called Neural Networks (NN). NNs also have several neural nodes that are connected and information is propagated between them. NNs have become increasingly popular due to its ability to learn complex patterns and relationships from data. This makes them suitable for a wide variety of tasks such as image and speech recognition, natural language processing, and even playing games.
I-D A Primer on Graph Neural Networks
In addition to the aforementioned learning techniques, there is a relatively new class of neural networks called Graph Neural Networks (GNN) [72, 73]. GNNs are based on graphs that are modeled based on two entities, namely nodes (ie., objects) and edges (ie., relationships between nodes). GNNs are well-suited for capturing complex relationships and dependencies in graph data, which is challenging for traditional neural networks. They can learn representations of nodes and edges based on their local neighborhood information. Second, GNNs can handle graphs of varying sizes and structures, making them applicable to a wide range of real-world problems. They can adapt to irregular and dynamic graphs while regular NNs fail. Third, pre-trained GNN models can be fine-tuned on specific graph-related tasks, enabling transfer learning. In this survey, we aim to provide a detailed study of GNN, discussing its design principles and architecture, state-of-the-art literature that covers how GNN has impacted various domains such as IoT, networking, spectrum awareness, and tactical systems, etc. Through this survey, we hope to make it easier for readers to learn and apply GNN to overcome some of the current research challenges of their respective domains of interest.
GNNs have been widely used to address a variety of network management issues such as resource allocation [74, 75, 76, 77, 78], spectrum sensing and sharing [79, 80], privacy [81], edge computing [82, 83], trajectory optimization [84, 85]. Flex-Net [74] uses GNN to jointly optimize the direction of communication and transmission power with low computational complexity as well as enable scalability and generalization capabilities. A GNN variant called Inductive Graph Neural Network Kriging (IGNNK) is used for secure cooperative sensing [79]. The advantages of IGNNK are that a trained GNN model can be transferred to a new graph structure as well as be used to generate data for virtual sensors for a given location as discussed in [86]. Decentralized management and control is one of the key features of next-generation wireless communication. However, privacy leakage is a major issue. To address this issue, in [81] the authors use GNN to design a privacy-preserving and privacy-guaranteed training algorithm for information exchanges among neighbors in wireless networks. In [82], the authors use a combination of GNN and Reinforcement learning (RL) to optimize the task offloading from IoT devices to mobile nodes that is robust to changes in network topology and mobile node failure. In [85], the authors present a GNN-enabled Advanced Actor-Critic (GNN-A2C) algorithm for aerial-edge IoT (EdgeIoT) system to jointly optimize trajectory and task offloading. With the continuous rise in the use of GNNs for a wide range of applications, it is important to keep track of existing literature works as well as have a clear idea of the challenges and opportunities provided by GNNs. Figure 2 shows our vision of GNN-enabled 6G networks with integrated ORAN architecture.
I-E Scope of the article
Survey of Surveys Detailed GNN GNN for GNN for Spectrum GNN for GNN for Tactical Research
Architecture IoT Awareness Networking Systems Directions
Ruiz et al. 2021[87]
Sato et al. 2020[88]
Yuan et al. 2023[89]
Zhou et al. 2023[90]
Abadal et al. 2022[91]
Wu et al. 2022[92]
Josephine et al. 2021[93]
Tam et al. 2024[94]
Dong et al. 2023[95]
Liu et al. 2021[96]
Jiang et al. 2022[97]
Gao et al. 2023[98]
Bilot et al. 2023[99]
Waikhom et al. 2021[100]
Tam et al. 2022[101]
Gupta et al. 2021[102]
Rahmani et al. 2023[103]
Suárez-Varel et al. 2023[104]
Munikoti et al. 2023[105]
Wu et al. 2021[106]
Zhou et al. 2022[107]
Longa et al. 2021[108]
He et al. 2021[109]
This Survey
TABLE I: Comparison of recent surveys related to GNN.
Figure 3: Overall Organization of the Survey.
Table I summarizes all the recent surveys related to GNN. Compared to the above works, in this survey, we focus on the following aspects:
  • Provide a detailed background about GNN and its architecture by discussing its components and functions. We also discuss the different types of GNNs and discuss related literature that uses different GNN types to solve a wide variety of problems. This is essential to provide the relatively newer reader with a great guide to start their journey.
  • Give a brief introduction to the Internet of Things and discuss the applications of GNN for IoT Network Intrusion Detection and IoT Data Fusion.
  • Provide an overview of spectrum awareness and elucidate the application of GNN for Radio Frequency Signal Classification and Radio Frequency Spectrum Sensing.
  • Delve into the basics of networking and discuss the applications of GNN for networking focusing on Network and Signal Characteristics Prediction, Routing Optimization, Congestion Control, Mobile Edge Computing, Digital Twin Networks, and Unmanned Aerial Networks.
  • Explore the basics of tactical systems and provide a qualitative discussion of the application of GNN for Tactical Communication and Sensing Systems and Localization
  • Deep-dive into the research opportunities and future research direction of GNN to motivate more exploration in the domains specific to next-generation wireless networks.
Notation Definition
N Number of Nodes in a graph
F Number of Features
F Feature Matrix
G Graph
V Vertices
E Edges
$i$ Index of node of interest
$j$ Index of neighboring node
$\textbf{x}_{i}$ Node Feature vector
$l$ Layer index
$L$ Total number of layers
$m_{i}^{(l+1)}$ Aggregated message of node $i$ at layer $l+1$
$\textbf{h}_{i}^{(l)},\textbf{h}_{j}^{(l)}$ Feature vectors of node $i$, node $j$ at layer $l$
$\textbf{h}_{i}^{(l+1)},\textbf{h}_{j}^{(l+1)}$ Feature vectors of node $i$, node $j$ at layer $l+1$
$\textbf{W}_{i}^{(l)},\textbf{W}_{j}^{(l)}$ Weight Matrices of node $i$, node $j$ at layer $l+1$
$\textit{N}_{i}$ Set of Neighbors of node $i$
$A(\cdot)$ Aggregator Function
$U(\cdot)$ Update Function
$\sigma(\cdot)$ Activation Function
$R(\cdot)$ Readout Function
$(\cdot)^{T}$ Transpose Operation
A Adjacency Matrix
D Degree Matrix
$e_{ij}$ Attention coefficient of node $i$ for neighbor $j$
$a(\cdot)$ Attention Mechanism Function
$\alpha_{ij}$ Normalized attention coefficient of
node $i$ for all neighbors $j\in N_{i}$
a Attention Mechanism Vector
||||| | Concatenation Operator
$\rm{inp}$ Input Vector
$\rm{enc}$ and $\rm{dec}$ Encoder and Decoder Output Vector
$f_{\rm{enc}}$ and $f_{\rm{dec}}$ Encoder and Decoder Functions
$\theta_{\rm{enc}}$ and $\theta_{\rm{dec}}$ Parameters associated with $f_{\rm{enc}}$ and $f_{\rm{dec}}$
$\mathcal{N}$ Number of elements in inp
$S_{\rm{inp}}$ and $S_{\rm{dec}}$ Sum of elements in $\rm{inp}$ and $\rm{dec}$
$||(\cdot)||^{2}$ Squared Euclidean Distance
$\epsilon$ Learnable or fixed scalar parameter
$f_{i}(\cdot)$ Graph Capsule Function
$\mathcal{F}_{i}$ Neighborhood feature value set
$\lambda_{ij}$ Edge weight between nodes $i$ and $j$
$x_{i}$ Node feature value
$p$ Number of instantiation parameters
TABLE II: List of notations and its definition.
Acronym Definition
6G Sixth-Generation
NextG Next-Generation
IoT Internet of Things
5G Fifth Generation
GHz Gigahertz
ML Machine Learning
GNN Graph Neural Networks
QoS Quality of Service
GRU Gated Recurrent Unit
ESN Echo State Networks
LSTM Long-Short Term Memory
CNN Convolutional Neural Networks
RL Reinforcement Learning
GCN Graph Convolutional Networks
GAT Graph Attention Networks
GraphSAGE Graph Sample and Aggregate
GAE Graph Autoencoders
GIN Graph Isomorphic Networks
RNN Recurrent Neural Network
GCapsNet Graph Capsule Networks
STGN Spatial-Temporal Graph Networks
MANET Mobile Ad Hoc Network
GCNN Graph Capsule Convolution
NID Network Intrusion Detection
RSSI Received Signal Strength Indicator
GTCN Graph-based Temporal Convolutional Network
DRL Deep Reinforcement Learning
MEC Mobile Edge Computing
DT Digital Twin
DDQN Deep Distributed Q-Network
UAV Unmanned Aerial Vehicle
UAN Unmanned Aerial Network
SAR Synthetic Aperture Radar
FL Federated Learning
TABLE III: List of commonly used acronyms and their definitions.
I-F Survey Organization.
The rest of the survey is organized as follows. We formally introduce GNN by discussing its inception and providing a detailed description of the architecture and types of GNN in Section II. In Section III, we discuss the applications of GNN for the Internet of Things, followed by applications of GNN for spectrum awareness, networking, and tactical systems in Section IV, Section V and Section VI, respectively. Finally, we provide our perspective on the opportunities that could motivate novel research directions utilizing GNN in Section VII before concluding in Section VIII. The overall survey overview is shown in Fig. 3
Figure 4: Examples of different GNN terminologies.
II Graph Neural Networks
In this section, we will provide a detailed introduction to Graph Neural Networks (GNN). To this end, we first present the commonly used terminologies in GNN (and neural networks in general) and their definitions in Section II-A, then we will focus on the architecture of GNN in Section II-B and then discuss different types of GNN variants in Section II-C.
II-A Terminologies and Definition
Before delving into the architecture and types of GNN, it is important to lay a strong foundation for basic terminologies used in Graph Neural Networks. Table II lists all important notations and their definitions. Table III lists all important acronyms and their definitions. Below is the list of important terminologies and their definitions used in GNN and the state-of-the-art literature discussed in later sections (Section III - Section VI).
  • Node. Nodes are the fundamental building blocks of a graph and represent an entity or an element in the graph.
  • Edge. Edges are the connections or relationships between nodes defining how nodes are related to each other.
  • Graph. A graph consists of nodes (blue solid circles in Fig. 4(a)) and edges (solid lines between two blue solid circles in Fig. 4(a)). In general, there are two types of graphs namely (i) Directed graphs and (ii) Undirected graphs as shown in Figs. 4(a) and (b), respectively. In a directed graph, edges have a specific direction, indicating a one-way relationship from one node to another. In an undirected graph, edges have no direction, indicating a two-way relationship where the connection is symmetric.
  • Neighborhood. Neighborhood refers to the set of nodes that are connected to the node of interest as shown in Fig. 4(b).
  • Node Feature. The attribute or properties associated with each node is called the node feature. These features include information about the node’s characteristics, attributes, or the context of the graph. Fig. 4(c) shows an example of node features where 0 and 1 represent the presence or absence of a feature for that node.
  • Edge Feature Similar to the node feature, the attributes or properties associated with each edge is called edge features. The edge feature captures the relationship between two nodes as shown in Fig. 4(d). In this example, the numbers represent the distance between the nodes.
  • Embedding. Embedding refers to the process of representing discrete variables as continuous, low-dimensional vectors. There are 3 types of embedding namely Node Embedding, Edge Embedding, and Graph Embedding.
  • Node Embedding. Node embedding involves learning a vector representation for each node in the graph such that nodes with similar network neighborhoods or connectivity patterns have similar embeddings. These embeddings capture the structural and relational information of nodes in the graph.
  • Edge Embedding. Edge embedding aims to learn vector representations for edges in the graph, capturing the relationship between connected nodes. Edge embeddings can encode information such as edge type, weight, or directionality.
  • Graph Embedding. Graph embedding involves learning a fixed-size vector representation for the entire graph. This representation summarizes the structural and topological properties of the graph, making it useful for tasks such as graph classification, clustering, or graph generation.
  • Aggregation. Aggregation refers to the process of collecting information at the node of interest from all the nodes in the neighborhood. Some of the standard aggregation techniques include Mean Aggregation, Max Aggregation, and Attention-based Aggregation.
  • Propagation. Propagation refers to the process of passing or propagating information within the graph. This process involves iteratively updating the node representations based on the information from their neighbors.
  • Activation Function. Activation Functions allow the defining of a nonlinear family of functions to capture the relationship between the input graph data and its representation. Some examples of non-linear activation functions include Sigmoid or Logistic, Hyperbolic Tangent, Rectified Linear Unit (ReLU), LeakyReLU, Parametric ReLU, Exponential Linear Units, Softmax, Scaled Exponential Linear Unit (SELU) and Gaussian Error Linear Unit (GELU), among others.
    Figure 5: Architecture of GNN depicting the three main stages.
  • Convolution. In traditional CNNs, convolution is referred as the process of sliding a filter (or kernel) over the input data while performing element-wise multiplication between the filter and the input and then summing up the results to produce a feature map. In GNN, convolution is referred as the process of updating the node features based on a transformation of aggregated information from the node’s neighbors.
  • Layer. In GNN (and NNs in general), a layer refers to a single computational step responsible for transforming the node features and propagating information across the graph. There are different types of layers such as input, hidden, readout, and broadcast layers. These layers can have specific functionality and are further divided into convolution layer, message passing layer, and pooling layer, among others.
  • Pooling. Pooling refers to the process of reducing the dimensionality of the graph while retaining important information. While pooling and embeddings have similar functionality, pooling works on a sequence of embeddings. For traditional CNNs, there are different types of pooling such as Max Pooling, Average Pooling, Global Average Pooling, Global Max Pooling, and Fractional Pooling. While for GNN, common pooling techniques include DiffPool [110], HIBPool [111], MinCutPool [112], NMFPool [113], Top-K Pool [114], SAGPool [115].
  • Feature Matrix. The feature matrix represents the 2-D representation of features of all the nodes in the graph. For example, if a graph has $N$ nodes, and each node has $F$ number of features, then the feature matrix F has a dimension of $N\times F$. Examples of node and edge feature matrices are shown in Figs. 4(e) and (f).
  • Adjacency Matrix. The adjacency matrix represents the connection between the nodes in a graph. An example of a Feature Matrix is shown in Fig. 4(g).
  • Degree Matrix. A degree matrix is a diagonal matrix containing information about the degree of each node. Here, the degree of the node indicates the number of terminating edges for a node. An example of a Feature Matrix is shown in Fig. 4(h).
  • Local Information. Local information refers to the information from either an individual node or information aggregated from the neighbors to compute a local graph representation.
  • Global Information. Global information refers to the information gathered by aggregated information from all the nodes in the graph to compute a global representation of the entire graph.
II-B Architecture of Graph Neural Networks
Graph Neural Networks (GNNs) are a class of neural networks designed to work with graph-structured data. Graphs consist of nodes and edges, where nodes represent entities (e.g., users in a social network or atoms in a molecule), and edges represent relationships or connections between these entities. GNNs are specifically designed to process and extract meaningful information from these graph structures, making them well-suited for a wide range of applications.
A simple GNN consists of the following three main stages as shown in Fig. 5.
  • Graph Representaion. A graph is represented as $G=(V,E)$, where $V$ is the set of nodes (vertices) and $E$ is the set of edges (connections) between nodes. Each node $i\in V$ typically has an associated feature vector $\textbf{x}_{i}$ that encodes information about the node. A GNN aims to learn representations of nodes that capture the graph’s structural and semantic information.
  • Message Passing. The core operation in GNNs is message passing, which enables nodes to exchange information with their neighbors in the graph. The message passing step is typically defined in two parts:
    • Aggregation Function: Each node aggregates information from its neighbors by computing an aggregation function (e.g., sum, mean, max) over the features of its neighbors. Mathematically, this can be represented as:
      $m_{i}^{(l+1)}=A(\textbf{h}_{j}^{(l)},\forall j\in N_{i}),$ (1)
      where, $m_{i}^{(l+1)}$ represents the aggregated message at node $i$ at layer $l+1$, $\textbf{h}_{j}$ represents the hidden state of neighbor node $j$ at layer $l$, $A(\cdot)$ represents the aggregation function, $N_{i}$ represents the set of neighbors of node $i$.
    • Update Function: Once messages are aggregated, each node updates its own representation based on the aggregated messages. This update function (U) can be implemented using neural networks (typically multi-layer perceptrons), and it can be represented as:
      $\textbf{h}_{i}^{(l+1)}$ $=U^{(l)}(\textbf{h}_{j},m_{i}^{(l+1)}),$ (2)
      $\textbf{h}_{i}^{(l+1)}$ $=\sigma(\textbf{W}_{i}^{(l)}\textbf{h}_{i}^{(l)}+\textbf{W}_{j}^{(l)}\textbf{h}_{j}^{(l)})$ (3)
      where, $\textbf{h}_{i}^{(l+1)}$ and $\textbf{h}_{i}^{(l+1)}$ feature vectors of node $i$ at layers $l$ and $l+1$, respectively, $\textbf{W}_{i}^{(l)}$ and $\textbf{W}_{j}^{(l)}$ are the weight matrices for nodes $i$ and $j$ at layers $l$. GNNs typically have multiple layers (or iterations) of message passing and updating. Each layer refines the node representations by considering information from a broader neighborhood in the graph.
  • Readout Function: After several message-passing layers, the final node representations are often aggregated to produce a graph-level representation or prediction. The readout phase computes a feature vector for each node $i$ (node-level) and is expressed as
    $y_{i}=R_{n}(\textbf{h}_{i}^{L^{T}}|i\in G)$ (4)
    Similarly, the readout phase can also compute the feature vector for the entire graph $G$ (graph-level) and is expressed as
    $y_{g}=R_{g}(A({h}_{i}^{L^{T}}|i\in G)),$ (5)
    where $R_{n}$ and $R_{g}$ are the readout function per-node and graph-level, respectively, $L$ is the total number of layers and $(\cdot)^{T}$ is the transpose operation.
II-C Types of Graph Neural Networks
In this section, we will discuss some of the most popular types of variations of GNN architectures namely Graph Convolutional Networks (GCNs), Graph Attention Networks (GAT), Graph Sample and Aggregated (GraphSAGE), Graph Autoencoders (GAE), Graph Isomorphic Networks (GIN), Graph Recurrent Neural Network (Graph-RNN), Graph Capsule Networks (GCaps) and Spatial-Temporal Graph Networks (STGNs).
Graph Convolutional Network. Graph Convolutional Networks (GCNs) are a type of deep learning model used for processing and analyzing data represented as graphs or networks. They were introduced to extend the power of CNNs to irregular data structures like graphs. CNNs are commonly used for tasks like image classification, object detection, image segmentation, and various computer vision tasks where spatial relationships are important while GCNs are used for tasks like node classification, link prediction, community detection, and other graph-based problems where understanding the relationships between data points is crucial. A GCN [116] consists of several convolutional layers and a single convolutional layer can be expressed as
[…]
where $\textbf{h}^{(l)}$ and $\textbf{h}^{(l+1)}$ are the feature vectors at layers $l$ and $l+1$, respectively. A is the adjacency matrix of the graph with self-connections, D is the degree matrix of A defined as a diagonal matrix where each diagonal element $D_{ii}$ is the sum of the corresponding row of A, $\textbf{W}^{(l)}$ is the weight matrix of layer $l$, and $\sigma$ is an activation function, such as ReLU or sigmoid. The applications of GCNs include Intrusion Detection [117], Cooperative Spectrum Sensing [118], Channel and Resource Allocation [119, 120], Modulation Recognition [121] and Joint Routing and Scheduling [122].
Graph Attention Network. The Graph Attention Network (GAT) [123] is a type of neural network architecture designed for processing data structured as graphs or networks. GAT is particularly useful for tasks such as node classification, link prediction, and graph classification, where information propagation and aggregation across graph nodes are crucial. GAT was introduced to improve how nodes in a graph aggregate information from their neighbors by applying attention mechanisms. Attention mechanisms are a crucial component in modern neural network architectures, and they have significantly improved the performance of various deep learning models across a wide range of tasks, especially in natural language processing and computer vision. At their core, attention mechanisms enable models to focus on specific parts of the input data when making predictions or generating output. The mathematical expression for an attention mechanism coefficient for GAT [123] can be expressed as
$e_{ij}^{(l)}=a(\textbf{W}\textbf{h}^{(l)}_{i},\textbf{W}\textbf{h}^{(l)}_{j})$ (7)
where $a$ is the shared attentional mechanism. This equation indicates the importance of node $j$’s feature to node $i$. To easily compare the coefficients across different nodes, the coefficient is normalized across all choices of $j$ using a softmax function and is given as
$\alpha_{ij}^{(l)}=\rm{softmax}(\textit{e}_{ij}^{(l)})=\frac{\rm{exp}(\textit{e}_{\textit{ij}}^{(l)})}{\sum_{\textit{j}\in N_{i}}\rm{exp}(\textit{e}_{\textit{ij}}^{(l)})}$ (8)
A LeakyReLU nonlinearity is applied to the Attention mechanism $a$ and can be expressed as
$e_{ij}^{(l)}=\rm{LeakyReLU}(\textbf{a}^{(l)^{T}}[\textbf{W}^{(l)}\textbf{h}_{i}^{(l)}||\textbf{W}^{(l)}\textbf{h}_{j}^{(l)}]$ (9)
where a is the weight vector for attention mechanism $a$, $(\cdot)^{T}$ and ||||| | are transpose and concatenation operators, respectively. Applying (9) in (8), we get
$\alpha_{ij}^{(l)}=\frac{\rm{exp}(\rm{LeakyReLU}(\textbf{a}^{(l)^{T}}[\textbf{W}^{(l)}\textbf{h}_{i}^{(l)}||\textbf{W}^{(l)}\textbf{h}_{j}^{(l)}])}{\sum_{j\in N_{i}}\rm{exp}(\rm{LeakyReLU}(\textbf{a}^{(l)^{T}}[\textbf{W}^{(l)}\textbf{h}_{i}^{(l)}||\textbf{W}^{(l)}\textbf{h}_{j}^{(l)}])}$ (10)
GATs are used in applications such as Channel Estimation [124], Malware Detection [125], Malicious Traffic Classification[126] Network Slicing Management [127], Task Offloading [128], Beamforming [129], Predictive Channel Modelling [130].
Graph Sample and AGgregatE (GraphSAGE). GraphSAGE is a graph representation learning technique that aims to generate embeddings for nodes in a graph. GraphSAGE starts by sampling a fixed-size neighborhood for each node in the graph. This neighborhood typically consists of the node itself and its immediate neighbors. The sampling process can be random or based on some specific strategy. For each node, GraphSAGE aggregates information from its sampled neighborhood based on an aggregator function. After aggregating information from the neighborhood, GraphSAGE generates a new embedding for each node. This embedding is learned through a neural network layer, which can be shared across all nodes or specific to each node. GraphSAGE combines techniques from graph sampling and aggregation to create node embeddings that capture both local and global graph structures. The mathematical modeling for GraphSAGE embedding generation [131] can be expressed as follows. First, each node $i$ aggregates the representations of its neighboring nodes into a single vector $\textbf{h}_{N(i)}^{l}$. Then, GraphSAGE concatenates the node’s current representation $\textbf{h}_{i}^{l}$ with the aggregates neighborhood vector $\textbf{h}_{N(i)}^{l}$ which is fed through a fully connected layer with non-linear activation function $\sigma$ to get $\textbf{h}_{i}^{l+1}$. This can be expressed as
$\textbf{h}_{N(i)}^{(l+1)}=A({\textbf{h}_{j}^{l},\forall j\in N_{i}}),$ (11)
$\textbf{h}_{i}^{(l+1)}=\sigma(\textbf{W}^{(l)}\cdot C(\textbf{h}_{i}^{(l)},\textbf{h}_{N_{i}}^{(l+1)}))$ (12)
where $N_{i}$ is the set of neighboring nodes of $i$, $\textbf{W}^{(l)}$ is the weight matrix of layer $l$, $A(\cdot)$ and $C(\cdot)$ are the aggregator and concatenator functions. There are three common types of aggregators namely Mean Aggregator, LSTM aggregator, and Pooling Aggregator. GraphSAGE is advantageous because it combines the power of both local and global information in graph data. GraphSAGE has been widely used for Intrusion Detection [132], MANET traffic analysis [133], Access Point Selection [134], Routing [135], Recommender Systems [136] and Network Traffic Prediction [137].
Graph Autoencoders. Graph Autoencoders (GAEs) are a type of neural network architecture designed for learning graph representations by encoding and decoding the graph structure. GAEs are used in unsupervised or self-supervised learning settings to capture meaningful latent representations of nodes or graphs. They enable the discovery of meaningful latent representations that can reveal the underlying structure and relationships within complex graphs. Researchers continue to explore variations and extensions of GAEs to improve their performance and applicability in different domains. GAE consists of two main components namely the Encoder which takes graph-structured data as input and maps it into a lower-dimensional latent space. It typically consists of graph convolutional layers that aggregate information from neighboring nodes while considering the graph’s topology. The output of the encoder is a compressed representation of the input graph and Decoder which takes the latent representation from the encoder and attempts to reconstruct the original graph. It maps the latent space back to the graph structure. The decoder’s goal is to generate a graph that is as close as possible to the input graph while minimizing reconstruction loss. The mathematical model for the encoder, decoder, and the loss functions [138] can be expressed as follows,
enc $=f_{\rm{enc}}(\rm{\textbf{inp}};\theta_{\rm{enc}}),$ (13)
dec $=f_{\rm{dec}}(\rm{\textbf{enc}};\theta_{\rm{dec}}),$ (14)
where enc and dec are the encoder and decoder output vector, $f_{\rm{enc}}$ and $f_{\rm{dec}}$ are encoder and decoder functions, inp is the input vector, $\theta_{\rm{enc}}$ and $\theta_{\rm{dec}}$ are parameters associated with $f_{\rm{enc}}$ and $f_{\rm{dec}}$, respectively. The reconstruction loss function is expressed as
$\rm{Loss}$ $=\frac{1}{\mathcal{N}}\sum_{i=0}^{\mathcal{N}}(||S_{\rm{inp},\textit{i}}-S_{\rm{dec},\textit{i}}||^{2})$ (15)
where $\mathcal{N}$ is the number of elements in inp, $S_{\rm{inp}}$ and $S_{\rm{dec}}$ are the sum of elements in inp and dec, respectively, $||(\cdot)||^{2}$denotes the squared Euclidean distance between the input and reconstructed features. GAEs are widely used for Anomaly Detection [139, 138], Wireless Sensor Network Topology Reconstruction [140], Indoor Localization [141], Intrusion Detection [142] and Task Assignment [143].
GNN Variant ML Library Used Related Literature
GCN Tensorflow [144] [145]
PyTorch [146]
GraphSAGE Tensorflow [147] [131]
PyTorch [148]
GAE Tensorflow [149] [150]
PyTorch [151]
GIN Tensorflow [152] [153]
PyTorch [154]
GCapsNet Tensorflow [155] [156]
PyTorch [157]
STGN Tensorflow [158] [159]
PyTorch [160]
TABLE IV: Example GitHub projects and related literature for different GNN variants.
Graph Isomorphic Networks. The Graph Isomorphism Network (GIN) is a type of neural network architecture designed for solving the graph isomorphism problem. The graph isomorphism problem involves determining whether two given graphs are isomorphic, meaning they have the same structure but might differ in the labels assigned to their nodes. A simple test called the Weisfeiler-Lehman test [161] can be conducted to determine if the graphs are isomorphic or not. The idea behind this test is to iteratively aggregate the labels of the nodes and their neighbors and hash the aggregated labels into unique new labels. If the labels of the nodes differ at some point in the iteration, then the graph is said to be non-isomorphic.
GINs are built upon GNNs and the key idea behind GINs is to aggregate information from neighboring nodes iteratively, updating the representation of each node based on its neighbors. By stacking multiple layers of such aggregation, GINs can capture increasingly complex graph patterns. Additionally, GINs incorporate techniques to ensure that they are permutation-invariant, meaning they can handle graphs with nodes in arbitrary order without affecting the result. A single layer of GIN [153] can mathematically expressed as
$\textbf{h}_{i}^{(l+1)}=\rm{MLP}^{(l+1)}((1+\epsilon^{(l+1)}h_{i}^{(l)}+\sum_{j\in N_{i}}\textbf{h}_{j}^{(l)}))$ (16)
where $\textbf{h}_{i}^{(l)}$ and $\textbf{h}_{i}^{(l+1)}$ are the representation of node $i$ at layers $l$ amd $l+1$, respectively. $N_{i}$ is the set of neighboring nodes of $i$, $\epsilon$ is a learnable or a fixed scalar parameter and $\rm{MLP}$ is a feed-forward multi-layer perceptron. GIN has shown promising results and has been applied to various tasks related to learning and classification [162, 163], network authentication [164], malicious node detection [165], network traffic prediction [166].
Graph Capsule Networks. Graph Capsule Networks (GCapsNet) are a type of neural network architecture inspired by capsule networks originally introduced for image analysis tasks. Graph Capsule Networks extend this concept to graph-structured data, allowing them to handle relational reasoning and hierarchical representations within graph data. In traditional CNNs, convolutional layers typically extract features from images by applying convolutional filters across spatial dimensions. Capsule networks introduce the idea of capsules, which are groups of neurons that represent instantiation parameters of entities within an image, such as object pose, scale, orientation, etc. These capsules allow for more robust and interpretable representations of objects compared to traditional CNNs. An example of GCapsNet called Graph Capsule Convolution Network (GCNN) is introduced in [167]. For node $i$ with node feature value of $x_{0}$ and a set of neighborhood node feature values $\mathcal{F}_{i}=\{x_{0},x_{1},x_{2},...,x_{j}\}$ including node $i$ the output of graph capsule function is $f:\mathbb{R}^{j}\rightarrow\mathbb{R}$ can be expressed as
$f_{i}(x_{0},x_{1},...,x_{j})=\frac{1}{|\mathcal{F}_{i}|}\sum_{j\in\mathcal{F}_{i}}\lambda_{ij}x_{j}$ (17)
where $\lambda_{ij}$ is the edge weight between nodes $i$ and $j$. For a graph capsule network that captures higher-order statistical moments, $f_{i}(x_{0},x_{1},...,x_{k})$ is replaced with a vector values capsule function $f:\mathbb{R}^{j}\rightarrow\mathbb{R}^{p}$, where $p$ is the number of instantiation parameters described in [168]. Then Equation (17) can be rewritten as
$f_{i}(x_{0},x_{1},...,x_{j})=\frac{1}{|\mathcal{F}_{i}|}\begin{bmatrix}\sum_{j\in\mathcal{F}_{i}}a_{ij}x_{j}\\ \sum_{j\in\mathcal{F}_{i}}a_{ij}x^{2}_{j}\\ \vdots\\ \sum_{j\in\mathcal{F}_{i}}a_{ij}x^{p}_{j}\end{bmatrix}$ (18)
where each row denotes statistical moments, such as mean, standard deviation, skewness, and other higher statistical moments. Due to the extensive features, GCaps finds its application in various domains, such as Recommendation systems [169], Classification [170], Malware Detection [171].
Figure 6: Common IoT Sensors and Intrusion Attacks.
Spatial-Temporal Graph Networks. Spatial-Temporal Graph Networks (STGNs) are a type of deep learning model designed to handle data that exhibit both spatial and temporal dependencies. They are particularly suited for tasks where understanding the interactions and dynamics across space and time is essential. STGNs operate on data represented as spatio-temporal graphs, where nodes represent spatial entities such as locations, sensors, and edges capture temporal dependencies such as timestamps, and time intervals.
STGNs integrate techniques from graph neural networks (GNNs) and temporal modeling to effectively capture spatial relationships and temporal dynamics within the data. They typically consist of multiple layers, including graph convolutional layers, temporal modeling layers, and attention mechanisms, to process and learn from the spatio-temporal graph data. STGNs are commonly used for Multi-object Tracking [172], IoT Intrusion Detection [173, 174] and Cellular Traffic Prediction [175, 176].
To further motivate the readers to use GNN in their research domain, Table IV lists some example GitHub projects as well as the related literature that can be referred to as a first step in understanding the design and implementation of different GNN variants.
III GNN for Internet of Things
Internet of Things refers to a network of interconnected devices that can communicate and share data with each other over the internet, without requiring direct human interaction. These devices are embedded with sensors, software, and other technologies that enable them to collect and exchange data, as well as perform specific functions. There are two major challenges with IoT. First, due to the vast number of devices connected through the internet, it is important to safeguard the devices from adversaries that may infiltrate the network and second, for applications that use different types of sensors to keep track of different entities, it is important to collect and process the data more efficiently. Towards this goal, IoT Network Intrusion and IoT Data Fusion have been extensively studied. We discuss the applications of GNN for IoT Network Intrusion Detection in Section III-A and IoT Data Fusion in Section III-B. Figure 6 shows representative IoT sensors and intrusion attacks.
III-A IoT Network Intrusion Detection
IoT Network Intrusion Detection (NID) refers to the process of monitoring and analyzing the traffic and behavior within an IoT network to identify and respond to unauthorized or malicious activities. IoT devices are increasingly being integrated into various industries and applications, ranging from smart homes and industrial automation to healthcare and transportation. However, the proliferation of IoT devices also presents new security challenges and vulnerabilities, making it crucial to have effective intrusion detection mechanisms in place for safeguarding IoT ecosystems, as these devices often have limited computing resources and may lack the robust security features found in traditional computing devices. By implementing effective intrusion detection systems, organizations can better protect their IoT networks from various threats, including malware, unauthorized access, data breaches, and other security risks. Below we discuss a few interesting approaches proposed for IoT NID. These are divided into general intrusion detection applications [132, 177, 178, 179] and Malicious logins and Malware detection [180, 181].
In [132], the authors present a GNN-based approach called Edge-Graph SAmple and aggreGatE aka E-GraphSAGE that allows capturing both the edge features as well as the topological information for network intrusion detection in IoT networks. Compared to traditional GraphSAGE [131] which mainly focuses on node features for node classification but falls short on edge classification using edge features. Classification based on edge features allows us to classify different network flows to the nodes into benign and attack flows which can be achieved using E-GraphSAGE. Similarly, in [177], the authors propose Anomal-E, a GNN-based approach to detect both anomaly and intrusion in computer networks using edge features and a graph topological structure in a self-supervised manner. In [178], the authors present a general framework for intrusion detection called Graph Intrusion Detection designed based on GNN for industrial IoT applications that are highly susceptible to cyber-attacks. This framework is designed to learn both the network topology as well as the GNN network parameter at the same time. While there are several ML-based intrusion detection techniques, they are not robust to common adversarial attacks. To address this challenge, in [179] the authors present a novel GNN-based model designed to process and learn from graph-structured information that keeps flow records and their relationships compared to other techniques that only focuses only on the different flows between the source and the destination hosts.
Another possible network intrusion scenario is malicious logins and the use of malware to gain access to the network. In [180], to address the challenge of malicious login the authors present a GNN-based solution called MLTracer. This framework consists of two key features. First, it enables differentiating crucial attributes of logins such as the source/destination IDs, roles, port, and authentication type, among others. Second, a GNN model in conjunction with CNN is used to detect malicious logins. In [181], the authors present a Network Flow GNN for malware detection and classification. The framework extracts the flow graphs to classify them using a novel edge feature-based graph neural network model. Three derived model variants are presented namely a graph classifier, a graph autoencoder, and a one-class graph neural network.
III-B IoT Data Fusion
IoT data fusion, also known as information fusion, sensor data fusion, or multi-sensor data fusion, refers to the process of integrating and combining data from multiple IoT devices and sensors to obtain a more comprehensive and accurate understanding of a physical or digital environment. The primary goal of IoT data fusion is to extract valuable insights, make informed decisions, and improve the overall efficiency and effectiveness of IoT systems. To implement IoT data fusion, various techniques and technologies can be used, including data preprocessing, sensor calibration, data alignment, feature extraction, ML algorithms [182, 183], and data visualization tools. The choice of fusion method depends on the specific use case and the characteristics of the data sources involved.
For example, in [182] the authors propose a deep learning method for fusing heterogeneous data collected from multiple monitoring stations using a Graph Convolutional Network specifically a spatiotemporal graph convolutional network. The overall idea behind this approach is to use the location of the stations and fuse the collected heterogeneous data and then predict the future trend based on global information rather than local information. Through this approach, better performance is guaranteed when compared to using data from individual stations. Applications such as surveillance systems and autonomous driving require the prediction of trajectories of multiple agents. These are affected by many factors such as each agent’s historical trajectory, and interactions between agents. In [183] the authors propose a trajectory prediction model based on the fusion of prior awareness and past trajectory information. Similar to [182], this paper also leverages the capabilities of a spatio-temporal convolutional network.
Figure 7: Applications of GNN for RF signal classification and prediction tasks.
Figure 8: Applications of GNN for RF spectrum sensing tasks.
IV GNN for Spectrum Awareness
As the number of wireless devices increases the spectrum becomes crowded where the same frequency band is shared among different services and applications. This causes interference and degradation in QoS which makes it critical to be aware of the spectrum to avoid interfering with licensed/unlicensed incumbent users and to enable effective use of the spectrum. To this end, many ML techniques have been used to enable spectrum awareness and GNN is one of the key enabling techniques due to its ability to model complex and dynamic environments. In this section, we shed light on the use of GNN for Radio Frequency Signal Classification (Section IV-A) and RF Spectrum Sensing (Section IV-B).
IV-A Radio Frequency Signal Classification
ML is known for its classification and prediction capabilities across domains. Traditionally, CNNs [184, 185, 186, 187, 188] and RNNs have dominated RF signal classification tasks such as modulation classification, signal type classification, RF Fingerprinting among others [188, 189, 190]. More recently, GNNs are also being leveraged for these tasks including but not limited to [191, 192]. Figure 7 shows common applications of GNN for RF signal classification and prediction tasks. Furthermore, GNN has also impacted other classification tasks such as anomaly classification [193] and traffic classification [194].
AvgNet [191] is an Adaptive Visibility Graph algorithm that maps time series data into graphs and can be used to classify various signal modulations. The main advantage of this algorithm is that it can automatically adapt into best-matching graphs, reducing the time consumed during the mapping process and aggregating both the local and global information of the signal sample. The proposed model is verified using simulation and the results show that AvgNet can effectively classify 11 modulations namely Amplitude Modulation - Double Sideband (AM-DSB), Amplitude Modulation - Single Sideband (AM-SSB), Phase Shift Keying (8PSK), Binary Phase Shift Keying (BPSK), Continuous Phase Frequency Shift Keying (CPFSK), Gaussian Frequency Shift Keying (GFSK), Pulse Amplitude Modulation (PAM4), Quadrature Amplitude Modulation (QAM16, QAM64), Quadrature Phase Shift Keying (QPSK) and Wideband Frequency Modulation (WBFM).
Another example of a signal modulation classification framework is DeepSIG [192] which is a hybrid heterogeneous modulation classification architecture designed based on the combination of Recurrent Neural Network (RNN), CNN and GNN. This framework can process radio signals with heterogeneous input formats such as in-phase (I) and quadrature (Q) samples, images mapped from IQ samples, and graphs converted from IQ samples. In addition to classifying the signals listed for [191], DeepSIG can also classify Minimum Shift Keying (MSK), Gaussian Minimum Shift Keying (GMSK), Orthogonal Frequency Division Multiplexing (OFDM), OFDM Quad-Carrier Quadrature Phase Shift Keying (OFDM-QPSK), Offset Quadrature Phase Shift Keying (OQPSK), OFDM Binary PSK (OFDM-BPSK), OFDM Quadrature Amplitude Modulation (OFDM-QAM16), Amplitude Shift Keying (ASK), Amplitude and PSK (APSK), Amplitude Modulation (AM), Frequency Modulation (FM), AM-MSK, FM-MSK, Pulse Amplitude Modulation (PAM), Continuous-Phase Modulation (CPM) and On-off key (OOK).
In addition to classifying the signal modulation, it is also important to detect and correct link failures and abnormal network behavior that may affect the large wireless network infrastructure. To address this problem, in [193], the authors present a new method that can detect wireless link anomalies based on GNN. Specifically, in this work, the authors use GIN architecture. The motivation for using GIN is due to the fact that the graphs of similar-looking anomalies have similar nodes and edges making these graphs isomorphic in nature. The effectiveness of the algorithm was verified using a real-world WiFi testbed.
GNN also finds its application for traffic classification scenarios where it is important to differentiate between traffic generated by various applications while identifying any encrypted or malicious traffic. To enable this, in [194], the authors present CGNN, a Chained GNN to effectively classify various real-world traffic scenarios and categorize them into the different applications as well as label them as ”normal”, ”encrypted” or ”malicious” traffic. Through extensive simulation, it is shown that compared to CNN and other state-of-the-art techniques, GNN, specifically CGNN outperforms them in terms of three key metrics namely precision, accuracy, and recall.
IV-B RF Spectrum Sensing
In the context of cognitive radios, spectrum sensing is performed to detect the presence of primary users (licensed users) in a given frequency band to enable opportunistic access for secondary users (unlicensed users). GNNs can analyze the received signals across different frequency bands and spatial locations, effectively capturing the spatial and spectral correlations among different spectrum channels. By learning the complex patterns in the received signals, GNNs can improve the accuracy of spectrum sensing, even in dynamic and heterogeneous environments. Figure 8 shows common applications of GNN for RF spectrum sensing tasks. In cooperative sensing scenarios [79] [118] [195], multiple sensors or nodes work together to gather and share information, improving the overall sensing performance beyond what individual sensors could achieve alone. This concept is prevalent in wireless communication networks, cognitive radio systems, distributed sensor networks, and other collaborative sensing applications.
In [118], the authors use a GCN-based cooperative spectrum sensing technique to address the hidden-node problem. The hidden node problem is primarily caused when a secondary user moves far away from the primary user or the fusion center thereby not participating in the decision-making process. Compared to traditional CNNs which assume the sensing matrices created by the fusion center as static but in reality, they are dynamically changing in the real environment. Thus needing a framework to feed dynamic graphs to the GCN for learning the features for accurate spectrum state classification. In [195], the authors investigate the use of GNN in conjunction with reinforcement learning techniques to optimize the energy efficiency of cooperative spectrum sensing wireless communication systems. The GNN is used to generate a feature vector from a distributed sensing node using graph-structured data and Q-learning based reinforcement learning is used to find the optimal solution.
V GNN for Networking
With the evolving number of wireless devices, the research community is working towards finding ways to improve the network performance to meet the quality of service requirements of the users. To this end, many ML-based solutions have been proposed in an effort to achieve the goal of better network performance. In this section, we will focus on applications of GNN for various networking applications such as Network and Signal Characteristics Prediction (Section V-A), Routing Optimization (Section V-B), Congestion Control (Section V-C), Mobile Edge Computing (Section V-D), Digital Twin Networks (Section V-E) and Unmanned Aerial Networks (Section V-F).
V-A Network and Signal Characteristics Prediction
Prediction applications include but not limited to Wireless LAN (WLAN) performance prediction [196, 197], Received Signal Strength Indicator (RSSI) prediction [198], network traffic prediction [199] and channel prediction [200]. In [196], the authors present a data-driven approach using GCN to predict the performance of next-generation WiFi networks. The designed model can predict the throughput of densely deployed WiFi nodes using Channel Bonding (CB). CB is a technique where two or more channels are combined thereby increasing the throughput between multiple wireless devices. In [197], the authors present HTNet, a specialized Heterogeneous temporal GNN that can extract features from dynamic WLAN deployment to predict its throughput. Through analysis of the unique graph structure of WLAN deployment graphs, HTNets can achieve the lowest prediction error in real-world experiments.
In [198] the authors use a combination of Random Dot Product Graphs (RDPG) and GNNs to address the problem of predicting wireless received signal strength (RSS) of wireless networks. RDPG is a spectral-based embedding method that provides a flexible framework for generating random graphs with various structural properties. Finally, in [199] the authors use a time-evolving graph to formulate the time-evolving nature of the user movements and propose a Graph-based Temporal Convolutional Network (GTCN) to predict the future traffic of each node in the wireless network. The GTCN can effectively learn intra- and inter-time spatial correlation between each node through the node aggregation method as well as can efficiently model the temporal dynamics of mobile traffic trends from different nodes through the temporal convolutional layer. In [200], the authors present an environment-information based channel prediction method to effectively predict the large-scale and small-scale fading parameters using GNN. First, scatterer-centered communication environment graphs is used to classify the scattering sources affecting the channel characteristics into effective, primary, and non-primary scatterers. The scatterers that produce multi-path and influence the channel characteristics are referred to as effective scatterers while the scatters that produce single propagation points are referred to as primary scatterers and all other scatterers are non-primary scatterers. After successful classification, the GNN-based channel prediction module accomplishes the task of predicting the large-scale and small-scale fading parameters using fully-connected propagation environment graphs and single-connected propagation environment graphs. Furthermore, a GNN-based multi-target regression model is designed for continuous channel parameters prediction.
V-B Routing Optimization
The key role of the network layer; the third layer of the Open Systems Interconnection (OSI) model is routing packets through the network. There are various routing protocols designed to facilitate the exchange of routing information among routers and enable them to make informed decisions about packet forwarding. The objective of routing protocols is to determine the best path or route for data packets from a source to a destination in a network based on predefined performance metrics (e.g. minimizing latency or minimizing energy consumption). Thus it is important to have an optimized routing algorithm that can improve efficiency by selecting the best routes, thereby reducing network congestion and improving overall network performance, resulting in better Resource Utilization.
Traditionally, techniques such as reinforcement learning, and recurrent neural networks were commonly used for routing optimization. Recently researchers have studied the applicability of GNN for routing optimization due to its ability to model and learn from graph-structured data. The key advantage of GNN compared to other methods is that by representing the network topology as a graph, where nodes represent routers or switches and edges represent links, GNNs can learn to predict the best paths for forwarding packets based on factors such as link capacities, congestion levels, latency and much more. The application of GNN for routing optimization has gained significant traction in the research community [122, 135, 201, 202, 203, 204, 205].
Figure 9: Workflow Diagram of GNN-enabled Mobile Edge Computing.
For example, in [122], the authors use a GCN-based DRL algorithm to jointly optimize traffic scheduling and routing. The key advantage of this algorithm is that there is no prior knowledge of the system, and all the information is obtained in run-time by interacting with the environment. In [135], they propose a reliable routing optimization algorithm using GraphSAGE. The algorithm consists of two components. First, a multi-path routing algorithm using breadth-first search method that can continuously predict the shortest path for different network topologies. Second is GraphSAGE-based routing performance optimization algorithm that creates network labels for training GraphSAGE and are used to learn graph features to predict the network performance for each path and select the best path for optimal network performance. This proposed method can minimize network congestion, improve network throughput, and reduce end-to-end delay and packet loss rate compared to traditional shortest-path routing algorithms.
RouteNet [201] is a novel network model-based GNN that can understand the complex relationships between topology, routing and input traffic and accurately predict the estimate of per-source/destination per-packet delay and packet loss. RouteNet can be generalized to arbitrary topologies, routing schemes, and traffic intensity thanks to GNN’s ability to learn and model graph-structured information. In [202] the authors explore the advantage of integrating GNN with Deep Reinforcement learning (DRL) to overcome the challenges of using DRL which are not well suited to learning from information structured as graphs and the limited generalizability of DRL. The proposed GNN-based DRL algorithm can learn and generalize to any given network topology. The proposed algorithm not only works for scenarios seen during the training process but also for network topologies that were not part of the training set. Similarly, in [203], the authors propose GRL-NET, a combination of GNN and DRL-based intelligent routing algorithm. Compared to using conventional methods to construct DRL agent, in this work, GNN is used due to its ability to learn from complex relationships among network topologies.
In [204], the authors present a GNN-based routing algorithm that leverages global information from a software-defined network to accurately predict the optimal path with the minimum average delay between the source and the destination node. In this work, GNN is primarily used because of its ability to generalize to different network topologies and varying data traffic scenarios. In [205], the authors propose a multi-path routing algorithm based on GNN to improve end-to-end transmission efficiency. First, a flow splitting scheme is designed to achieve multi-path transmission, and second, GNN is used to predict the link delay and select the most-efficient efficient path. The proposed algorithm outperforms the existing solutions in terms of time overhead, end-to-end delay, throughput, and generalizability to different topologies.
V-C Congestion control
Congestion control refers to the techniques and mechanisms used in computer networks to regulate the flow of data packets, preventing network congestion, and ensuring efficient data transmission. When network resources such as bandwidth, buffers, or processing capacity are overwhelmed with data, congestion occurs, leading to packet loss, increased latency, and reduced throughput. Congestion control aims to manage this situation to maintain network stability and performance. While traditional ML algorithms are extensively used for congestion control, now the research community is investigating the use of GNN to efficiently control the congestion in small and large-scale networks. Some of the research work includes CongestionNet [206], CONAIR [207] and RouteNet-Fermi [208].
In [206], the authors present CongestionNet, a graph-based deep learning method using GAT to quickly predict routing congestion hotspots. GAT was used because the GAT model allows the network to make higher frequency decisions compared to other methods and also performs better on inductive inference problems where there are no labels during the inference time. CONAIR [207] is a proactive Congestion Aware Intent-Based Routing architecture that allows selection of best communication link based on quality of service (QoS) using GNN. GNN allows the prediction of the future state of the network based on QoS metrics at each input node within the network. Furthermore, this approach can be applied to dynamic networking topologies with models trained on different scenarios. In [208], the authors introduce RouteNet-Fermi, a custom GNN model that can be used to accurately predict the delay, jitter, and packet loss in the network. One key advantage of this method is that it can also provide accurate estimates in networks that were not in the training set.
V-D Mobile Edge Computing
Mobile Edge Computing (MEC) is a decentralized computing paradigm that brings computation and data storage closer to the location where it is needed, rather than relying solely on centralized cloud data centers. In edge computing, data is processed and analyzed locally on edge devices or servers, often at or near the data source, such as IoT devices, sensors, and other endpoints. Edge computing and GNNs can be combined to create intelligent, real-time decision-making systems at the edge of the network. This combination is particularly powerful for applications that involve analyzing data from a network of interconnected devices or sensors, where GNNs can capture complex relationships and patterns in the data. Figure 9 shows the workflow of GNN-enabled MEC. The workflow can be divided into six steps. First, the user equipment (such as a cellular device or cellular-enabled UAV) sends the computation/task request and relevant information (CTRI) to the base station. Second, the base stations aggregate all the task information and QoS requirements and forwards the combined combined CTRI to the Edge Server Controller (ESC). Third, The ESC collects the information, generates the graph representation and the GNN decision agent (usually coupled with an RL) decides the best edge server capable of executing the task that satisfies the user QoS and sends a computation/task allocation (CTA) message to the edge server. Fourth, The BS then relays this CTA message to the user. Fifth, the user sends the task data to the edge server through the BS. Sixth, the edge server executes the task and sends the computation/task output (CTO) back to the user. Each of these steps is clearly labeled with numbered blue circles for cellular-enabled UAVs, numbered red circles for cellular users, and numbered purple circles for the combined messages.
In general, MEC can be divided into two parts. The first is Secure and Cost-efficient Computation [209, 210, 211] and the second is task offloading [212, 213, 214, 215, 216]. In Section III-B, we discussed the use of data from different sensors for prediction tasks. These data are typically processed in a centralized server or a group of edge-servers. The final data obtained may be malicious when the server is under attack by an adversary, or the edge server may be malfunctioning due to other factors. In order to correctly recover the final output in either of these situations, it is important to introduce additional security measures. In [209], the authors propose the use of a generalized graph NN that would help identify the edge server types and categorize them into faithful and malicious. This approach has lower complexity and faster convergence than conventional neural networks. In [210], the authors build a novel modeling framework for optimizing system cost based on distributed GNN processing over heterogeneous edge servers. They also developed an incremental graph layout strategy to address the dynamic evolution of GNN’s input data graph and an adaptive scheduling algorithm to balance the tradeoff between overhead and system performance. HGNAS [211], is the first Hardware-aware Graph Neural Architecture Search framework for resource-constrained edge devices. The key feature of HGNAS is that it can automatically search for optimal GNN architectures that maximize given performance metrics such as task accuracy or computation efficiency.
Task offloading in MEC refers to the process of transferring computational tasks from mobile devices (such as smartphones, tablets, or IoT devices) to nearby edge servers or nodes located at the edge of the network infrastructure. In traditional cloud computing models, all computation is typically performed in centralized data centers, which can result in high latency and network congestion, especially for latency-sensitive applications running on mobile devices. Task offloading in MEC aims to alleviate these issues by bringing computation closer to the end-users, reducing latency, and improving the overall quality of service. GNN and Task offloading have been extensively studied and are applied to various applications [212, 213, 214, 215, 216]. For example, in [212] the authors propose a task offloading mechanism based on graph neural network called graph reinforcement learning-based offloading (GRLO) framework is paired with an actor-critic network to overcome the challenges of Heuristic algorithms and deep reinforcement learning-based approaches that heavily depend on accurate mathematical models for the MEC system, and deep reinforcement learning techniques does not make fair use of the relationship between devices in the network. Similarly, in [213], the authors propose a meta-reinforcement learning task offloading algorithm called GASTO to overcome the challenges associated with the generalization and robustness of offloading algorithms in dynamic and uncertainties in the real-world environment. In [214], the authors propose a task offloading mechanism based on GCN which generates the graph and analyzes the integrated relationship from features such as task characteristics, network conditions, and available resources at the edge. GCN enables improved decision-making, enhanced resource utilization, and optimized performance in the edge network. In [215] the authors present an end-to-end fine-grained computing offloading model with the goal of achieving better load balancing. To this end, a deep graph matching method based on GNN is used to select the best node for offloading and can be applied for dynamic and large-scale networks and achieve faster execution speed. In comparison to existing methods for task offloading and load balancing, the proposed method can reduce the network load imbalance efficiently and due to its shorter execution time, the proposed scheme is suitable for real-time scenarios with multiple subtasks and edge nodes. Similarly, in [216], the authors propose a GCN-RL based algorithm for Container Scheduling for NextG IoT user task offloading and resource scheduling architecture with the objective of minimizing average task completion time for fine-grained and dynamic container scheduling problems.
Figure 10: Workflow diagram of GNN-enabled DT networks.
V-E Digital Twin Networks
One of the most promising enabling techniques for future IoT and next-generation networks is Digital Twins (DT) [217]. Digital Twins are a virtual replica of the real world which enables researchers to create an accurate model of the real-world environment to test and improve their ML algorithms before integrating with real-world hardware. Figure 10 shows an example workflow diagram of GNN-enabled DT network.
Over the past few years, DT has gained significant interest from the research community spanning across different domains. GNN is envisioned to be a powerful framework for analyzing and extracting information from complex network data [218, 219]. In addition to traditional optimization and ML algorithms, GNN has been extensively used together with DT in applications such as network optimization [220, 221, 222, 223], 6G Edge Networks [224] and resource allocation [225].
For example, in [220], the authors present a scalable DT of 5G network to model the interactions between multiple network slices into a graph and use a GNN-based virtual representation model that captures the hidden interdependencies among nodes based on the aggregation of information from multiple neighbors. Here, the DT is used to explore the complex traffic generated by the network slices and accurately predict the end-to-end slice latency. The DT designed based on GNN is shown to have higher accuracy in estimating the end-to-end latency for different topologies, traffic distributions, and QoS constraints that were not part of the training process. Similarly, in [221], the authors propose a DT-enabled deep distributed Q-network (DDQN) framework that constructs a digital replica of physical slicing-enabled beyond 5G networks that can simulate complex environment and predict its dynamic characteristics. Here, the DT is represented using graph and GNN to learn the complicated relationships of the network slice and the network states are forwarded to the DDQN agent to learn the optimal network slicing policy. It is shown that the proposed DT model achieves optimal and autonomous network management of complex slicing-enabled beyond 5G networks.
Figure 11: Typical components of UANs.
Figure 12: Applications of GNN for UAN tasks.
Figure 13: Application of GNN for tactical system tasks.
In [222], the authors propose TwinNet, a GNN-based DT model that can take into account the complex relationship between different service level agreements metrics such as queuing policy, network topology, routing configuration, and input traffic. It is shown that GNN-enabled TwinNet can generalize to its input parameters, topologies, routing and queuing configuration, and can accurately predict the end-to-end path delays in 106 unseen real-world topologies from Internet Topology Zoo [226]. Flow-aware Digital Twin (FlowDT) [223], is a GNN-based framework to model computer networks, where flow refers to the set of packets that share common characteristics such as the same protocols (UDP/TCP), IP addresses or port numbers of source and destination node. FlowDT supports the creation and destruction of flows and also provide accurate estimates of per-flow metrics such as delay and jitter.
In [224], the authors analyze the self-healing mechanism in 6G edge networks using DT. Self-healing 6G networks refer to the next-generation of wireless communication infrastructure that possesses advanced capabilities to automatically detect, diagnose, and mitigate issues or disruptions within the network without human intervention. These networks are designed to be highly resilient and reliable, capable of maintaining connectivity even in challenging conditions or in the face of malicious attacks. GNN is used to design a performance prediction mechanism that can accurately predict the network performance and detect any abnormal links or nodes. In [225], the authors present a DT architecture for Terahertz (THz) wireless network and represent the DT network as a graph. The generated graph is used in a GNN-based multi-set distributed message propagation algorithm to solve power allocation and user association for THz networks.
V-F Unmanned Aerial Networks
Unmanned Aerial Networks (UANs) refer to networks formed by interconnected Unmanned Aerial Vehicles (UAVs), also known as drones, operating cooperatively to achieve various objectives. Due to the portable nature of the UAVs, they can be dynamically deployed when required and play a crucial role in establishing temporary connectivity in the event of disaster or infrastructure failure and enable high data rate network connectivity. Figures 11 and 12 show the Typical components of UANs and applications of GNN for UAN tasks, respectively. ML techniques such as RNNs and RL are commonly used for UAN problems such as spectrum access and flight control for microwave, millimeterwave, and Terahertz frequency bands [227, 228, 229, 230, 231, 232], UAV swarm collaborative beamforming [233, 234, 235, 236]. In addition to traditional ML techniques GNNs are also commonly used for UANs problems such as Sensing [237] and networking applications such as trajectory planning [238, 239], resilient UAV swarm communication [240], Flying Relay Location and Routing Optimization [241].
In [237], the authors propose a GCN and GAT-based LSTM model to accurately predict the real-time and future Air Quality Index (AQI) using aerial-ground sensing framework. This model uses a lightweight Dense-MobileNet model [242] to achieve energy-efficient end-to-end learning from features extracted from images captured by the UAV. This framework allows for federation across various institutions to collaboratively learn a well-trained global model to accurately and securely monitor the AQI and also expanding the scope of UAV swarm monitoring. In [238], the authors present a GCN-based framework to solve the cooperative trajectory design problem formulated based on modeling the relationship between the ground terminal and UAV-base stations (UAV-BS). This framework enables UAV-BS to efficiently manage the time-varying local observation such as current location, average throughput, and scheduling and facilitate cooperation between UAV-BSs through information exchange. Due to the lack of global information and a globally optimal solution, a Q-learning-based multi-agent RL algorithm is used by the UAV-BS to learn a distributed trajectory policy which is modeled as a Markov game. Here, the message passing by the GCN framework ensures cooperative behavior among UAV-BS and significantly enhances the network’s performance. In [239], the authors propose a multi-dimensional trajectory prediction model for UAV swarms by integrating GNN with a dynamic graph neural network (DynGN). The DynGN uses an encode-decoder structure with a GCN to process the dynamic adjacency matrix, extract spatial features, and gated recurrent unit to model the temporal dynamics and capture the evolution of UAV interactions over time. This model processes both the dynamically changing network configuration and the trajectory data of the UAV swarm simultaneously.
In [240], the authors study the self-healing problem of UAV swarm network to re-establish network connectivity in the event of unpredictable destruction. These destructions may affect one or more UAVs simultaneously and can either be one-off events or recurring events. To address the one-off events, a GCN-based framework is designed to find the recovery topology of the UAV swarm network in a timely manner. This framework is then extended by incorporating a monitoring mechanism to detect the UAV destruction and a self-healing trajectory planning algorithm based on GCN to address the recurring events of destruction. Finally, a meta-learning scheme is designed for the GCN to reduce the execution time complexity and ensure faster convergence of GCN. In [241], the authors propose a GNN-based framework for the optimization of UAV locations and relay paths in UAV-relayed 6G IoT networks. A two-stage GNN model is developed to decouple the problem of optimal relay path selection and UAV position optimization and is solved by a separate GNN model. Here, both stages are trained using reinforcement and unsupervised learning with no training data or prior knowledge of optimal solutions. First, the locations of the UAVs are optimized and second, the best relay path is determined based on the optimized location of the UAVs from the first step. Results show that the proposed model performs better than brute search with lower time complexity, scalable to very large networks with dynamically changing environments.
VI GNN for Tactical Systems
Tactical systems play a crucial role in modern warfare by enhancing the effectiveness, survivability, of military forces operating in dynamic and challenging environments. It is important to maintain a safe communication system and to identify and localize any imminent threat. To this end, in this section, we discuss the application of GNN for Tactical Communication and Sensing Systems (Section VI-A), Target Recognition (Section VI-B) and Localization (Section VI-C). Figure 13 shows applications of GNN for tactical system tasks.
VI-A Tactical Communication and Sensing Systems
Tactical Communication and Sensing Systems refer to communication networks and technologies specifically designed for military or tactical operations. These systems aim to provide reliable, secure, and efficient communication among various military assets, including soldiers, vehicles, sensors, and other connected devices, in dynamic and challenging environments. GNNs can be used to enhance situational awareness on the battlefield. By modeling the battlefield as a graph where nodes represent military assets (e.g., vehicles, soldiers) and edges represent relationships or interactions (e.g., proximity, communication), GNNs can analyze the evolving battlefield situation in real-time. This can aid in decision-making and mission planning [243, 244, 245, 246, 247], training simulations [248], human-intelligence collaboration [249], airspace complexity prediction [250] and tactical mobile networking [251].
In [243], the authors propose context-surrounding graph neural networks with numbers (CS-GNN-N) for knowledge-based reasoning for military decision support. One of the major components is GNN-based learning which utilizes two smoothness metrics, namely feature smoothness and label smoothness to measure the quantity and quality of neighborhood information of nodes. In [244], the authors present a GNN-based mission planning approach for unmanned ground vehicles (UGV) to generate optimal mission planning results using recursive graph grammar to describe the task assignment process of multiple UGVs as an irregular graph structure and by coupling both task assignment as well as path planning. The performance of the proposed algorithm was verified both using simulation and real-world UGV robot. In [245], the authors propose GraphNearbyAlert (GNA) framework to effectively broadcast cooperative awareness messages to a group of vehicles. The results show that it is sufficient to transmit very little information about the topology of the graph to extract information from the shared data and can be used for decision making in cooperative transportation systems while preserving privacy. In [246] the authors present a GCN framework integrated with deep Q network based RL-based to achieve an efficient and reliable multi-agent decision making system for safe and efficient operation of connected autonomous vehicles in intelligent transportation systems. The GCN generates the features of the network based on the topology and is used by the RL to generate optimal decisions. In [247], the authors present an improved single-agent GNN-based algorithm for decision making in autonomous vehicles, where the GNN is first used to generate features for a multi-agent system and then transferred to a single-agent system. The advantage of this approach is that the newly trained single-agent model can then be used to train scenarios with multiple agents. This design significantly reduces the time complexity of training a multi-agent model and improves the convergence speed of the model.
In [248], the authors introduce a novel GNN based behavior prediction model to train and develop synthetic characters for roles in the Rapid Integration and Development Environment (RIDE). RIDE is a unity-based military training simulation environment that facilitates rapid development and prototyping of simulated environments in direct service of the Army and other Department of Defense entities. Through simulation experiments, the authors show that a GNN model trained using pseudo-human behavior can be used as a behavior predictor and help with better cooperation between human agents. In [249], the authors propose a GNN-based DRL framework to autonomously accomplish real-time resource dispatching and allow human intervention to accomplish complex tactics. In this design, the GNN is used to generate an integrated scheduling model of detection, tracking and interception and transformed into a sequential decision problem. The GNN model is integrated with the proximal policy optimization (PPO) algorithm to learn the air defense environment. This framework enables human-intelligence collaborative dynamic scheduling for emergency response scenarios. In [250], the authors propose a multi-modal adaptive STGN framework to simultaneously explore spatio-temporal dependencies in the airspace network with the objective to effectively learn the diverse spatial relationship and adaptively adjust the impact of different spatial models. The results show that the framework can harness differing spatial models and achieve high generalization performance across different temporal patterns compared to state-of-the-art methods. In [251], the authors present a GCN-based multi-agent DRL routing algorithm for multi-sink tactical mobile sensor networks to overcome link-layer jamming attacks. The proposed algorithm captures the hop count to the nearest sink (i.e., ), the one-hop delay, the next hop’s packet loss, and the energy cost of packet forwarding. The results show that the proposed scheme outperforms compared to other state-of-the-art algorithms.
VI-B Target Recognition
In addition to its use for decision-making, GNN has widely been used for automatic target recognition. Synthetic Aperture Radar (SAR) images are commonly utilized in military applications for automatic target recognition (ATR). SAR is a remote sensing technology that uses radar to create high-resolution images of the Earth’s surface. SAR works by emitting radio waves at a target and then measuring the reflection of those waves. Unlike optical imaging, SAR can operate day or night and in various weather conditions because it doesn’t rely on visible light. Due to this ability, SAR images are widely used in the military for ATR [252, 253, 254]. In addition to SAR images there are also other techniques used for target recognition such as Hyperspectral images [255, 256] and Infrared thermal imaging [257]. In [252], authors use GNN to identify ground-based objects such as battle tanks, personnel carriers, and missile launchers, among others. By identifying the type of objects, it would allow the military to determine if the objects are friendly or adversarial. However, the classification is susceptible to the noise level in the SAR images which may degrade the performance of the GNN classifier. The effect of an adversarial attack on GNN-based SAR image classifier has been extensively in [253]. Furthermore, extraction of features from the original SAR image requires a large amount of data and is time-consuming. To overcome this issue, in [254], the authors use SAR image target pixel gray-scale decline by a graph representation. Here, the gray-scale SAR image is divided into subintervals and a node is assigned to represent each pixel with declined order of pixel gray-scale in the subinterval. The raw SAR image is then transformed to construct a graph structure which is used by the GCN to extract the features of the graph structured data and then classify the target.
Hyperspectral imaging (HSI) is a process of capturing images with a wide spectrum of light across many narrow wavelengths. Due to this, compared to traditional imaging, HSI can help extract more detailed and accurate information such as precise color and material identification. In [255], the authors propose CFGC, a cognitive fusion of GNN and CNN for enhanced hyperspectral target detection (HTD), where CNN is used to extract multi-scale features and then a depth-wise separable convolution-self attention (DSC-SA) mechanism is used to reduce the number of CNN parameters while enhancing the feature space. The GNN local aggregation and global attention block (LAGAB) [258] mechanism to extract the global information. Compared to state-of-the-art detection algorithms results show that CFGC achieves high robustness and better detection performance. Similarly, in [256], the authors propose a spectral graph transformer-based HTD method where a pixel spectrum is first constructed as a spectral graph and a GCN is used to extract the local information of the spectrum. Then, a self-attention transformer is used to learn the global information of the spectrum/ The framework also integrates a graph contrast clustering algorithm to enable high spectral discrimination ability. In [257], the authors develop a GCN-based adaptive graph reasoning method for aerial target pose estimation modeled based on the eye-tracking properties of humans. The GCN is employed to derive the dynamic prediction of the object motion based on the spatio-temporal and visual data. The GCN uses the historical pose trajectories to predict the corresponding pose in the next time instant. This framework can help detect the target as well as its trajectory.
VI-C Localization
Localization refers to the process of determining the location or position of an object or entity within a given environment. It is a fundamental task in various fields, including robotics, navigation, mobile computing, and wireless communication. Accurate localization is crucial for applications such as GPS-Based Navigation, Augmented Reality (AR), and autonomous vehicles. GNNs are used for localization in certain contexts due to their ability to capture spatial relationships and dependencies among different elements in a graph-based representation of the environment. GNNs can integrate information from different types of sensors or data sources within the same graph structure. This is beneficial in scenarios where multiple sensors (e.g., GPS, Wi-Fi, LIDAR, cameras) are available for localization. Localization data can be noisy and uncertain due to factors like signal interference, multi-path effects, and environmental changes. GNNs can learn to handle such noise and uncertainty by aggregating information from neighboring nodes and making robust predictions.
GNNs are widely used for indoor localization [259] [260], large-scale network localization [261] and false data injection attack localization [262]. For example, in [259] graphs are generated using the RSSI from Wi-Fi access points to localize the device into certain predefined areas using GCN. Similarly, in [260], Zhang et al. proposed a novel WiFi domain adversarial graph convolutional network model for indoor localization using RSSIs between waypoints and WiFi access points. The difference in the two works is that in the latter unlabeled data and multiple data domains are considered and a semi-supervised domain adversarial training scheme was used to efficiently use the unlabeled data as well as different data distributions across data domains. While GNNs are widely used for indoor localization, they also find application in large-scale network localization such as [261] where a network with 500 nodes was considered in comparison to 25 nodes/devices in [259]. In [262], the authors propose a GNN-based approach to identify the presence of a false data injection attack as well as localize the source of the attack.
VII Opportunities and Roadmap ahead
As seen in the article, GNNs have been widely adopted for a wide range of applications but several areas are in search of breakthroughs to make GNNs more impactful. This provides an opportunity for the research community to further develop new features for GNN as well as apply GNN for even broader research areas. To this end, in this section, we discuss the possible avenues of research to improve the capabilities of GNN and research areas where GNN will play a crucial role in IoT and NextG networks. Figure 14 shows the key areas of research opportunities.
  1. 1.
    Scalable and Reliable GNN. Scalability is one of the crucial aspects to gauge the success of any ML algorithm and GNN is not an exception. As of now, most of the GNN algorithms are tested on a smaller network topology and present good results with a slightly larger network to show the scalability. However, with the evolving number of IoT devices and wireless devices, it is crucial to develop algorithms that can scale up to thousands or hundreds of thousands of devices without any possible degradation in performance. GNN promises to be a great tool that can enable scalable ML models, it is important for the research community to further investigate and rigorously test their algorithm from the perspective of scalability on real networks.
    Reliability is another important area that requires further attention. As of now, most of the GNN-based algorithms designed assume a perfect environment with pre-defined interference, channel, and other relevant models for system design. It is important to design GNNs that are robust to adversarial attacks, sensor malfunction, data corruption, and other unforeseen conditions. This is especially important when there are millions of connected devices and data from one or more sensors should not impact the whole network. This can be addressed by systematically identifying, quantifying, and rectifying anomalies or security risks. This promises to be an interesting area of research and will significantly benefit in maturation and adoption of GNN across different domains.
  2. 2.
    Capturing Network Complexity. GNNs are often trained on a smaller network size and relatively smaller dataset which are insufficient to capture the underlying complexity of the network. Another possible situation with collected datasets is that they can be imbalanced. This imbalance can occur when more data points are collected for a perfect scenario without any abnormal or noisy cases. When this data is used for training a GNN model, the developed model will work well for perfect cases while failing in cases with noise or any perturbations. Techniques such as Data Augmentation [263], and Generative Adversarial Network [264] can be integrated with GNN as one of the possible solutions to address the challenge of limited data. The research community is also encouraged to explore the integration of Transfer Learning, Meta Learning as additional possible solutions.
    Figure 14: Overview of Research Opportunities.
  3. 3.
    Interpretable and Explainable GNN. One of the major challenges with any ML-based algorithms is the lack of interpretability. Interpretability can be defined as the measure of how much the user can understand the decision of the model. In other words, if the user can understand and predict the output of the ML algorithm then the ML algorithm/model is said to be highly interpretable. Similarly, the users should also be able to understand the factors that affect the performance of the GNN model and ensure that the decisions align with human intuition and domain knowledge. However, in the wireless domain (specifically IoT), with more nodes and heterogeneous data, it is challenging to design a highly interpretable and explainable GNN model. A good starting point for research along the design of interpretable and explainable GNN is GNNExplainer [265], the first-general, model-agnostic approach that can provide interpretable explanations for any GNN-based model on any graph-based learning tasks. This line of research is very important to further enhance GNN’s features, making it a reliable and preferred ML technique across many domains.
  4. 4.
    Preservation of Privacy and Security. The current applications of GNN for IoT and wireless networks collect data from different sensors and users as well as data shared between sensors (or users). While this is not an area of concern for simulation-based implementations for proof-of-concepts implementations, it is important to integrate privacy and security measures into the GNN algorithm. One of the most promising and well know techniques for preserving privacy and security is Federated Learning (FL). In FL, instead of sharing the data between nodes/sensors, only the model weights are shared. This enables collaborative learning without compromising data privacy and security. An example of GNN using FL is FedGraphNN[266] which is an open FL system for research on Federated GNNs. The integration of GNN with Federated Learning is an interesting and important area of research for future IoT and next-generation wireless networks.
  5. 5.
    Zero-Touch GNN. In the current implementation of GNNs, the algorithm designer has to grapple with the design of GNN making the process tedious for complex network scenarios. Automated Machine Learning (AutoML) [267] promises to be a solution to overcome the tedious design process. AutoML is an enabling technique to improve the implementation of ML models such as GNN by automatically learning the important steps of the ML model design process. There is some progress towards Zero-Touch GNN such as Auto-GNN [268] which helps in finding the optimal GNN architecture within a predefined search space and Auto-STGCN [269] which helps in generating optimal STGCN models automatically by quickly scanning over the parameter search space. Zero-Touch GNN will play an important role in increasing the efficiency of the GNN modeling process and requires significant attention from the research community to make it a widely adopted strategy in designing the GNN model for the future of wireless networks.
  6. 6.
    Bridging the Sim2Real Gap. In most of the existing literature, the researchers tend to focus more on two aspects of validating the performance of designed the GNN model. First is the use of open-source datasets and second is the simulation environment. This creates a gap between the performance of the proposed GNN algorithm in a simulation environment and the real-world environment. There are several National Science Foundation (NSF) Platforms for Advanced Wireless Research (PAWR) platforms such as Colosseum [270], Platform for Open Wireless Data-driven Experimental Research (POWDER) [271], Cloud Enhanced Open Software Defined Mobile Wireless Testbed for City-Scale Deployment (COSMOS) [272], Agriculture and Rural Communities (ARA) [273], Aerial Experimentation and Research Platform for Advanced Wireless (AERPAW) [274] and other testbeds such as ARENA [275], CloudRAFT [276], UnionLabs [277]. All these platforms are available to the research community to conduct real-world experiments. The researchers can share the test scenario, code, and dataset with the community to accelerate the growth of GNN as well as next-generation wireless networks.
  7. 7.
    Cross-Layer Optimization. In wireless networks, each layer is interconnected, and changes made in one layer have direct or indirect effect on other layers. Currently, most of the state-of-the-art literature of GNN focuses on optimizing only one layer, for example, the physical layer sensing or network layer routing. However, it is important to jointly optimize more than one layer to achieve better network performance [278, 279]. A rare example of GNN-based cross-layer optimization is GCLR [280] which is a GNN-based cross-layer optimization system for Multipath TCP protocol. This framework helps in predicting the expected throughput and the best routing path based on the given network topology and multipath route candidate set. While traditional optimization has been extensively used for cross-layer optimization problems [281, 282, 283, 284, 285, 286], GNN promises to be one of the best solutions for large-scale, dynamically changing networks.
  8. 8.
    Zero-Trust Networking. Security is always an active research area for IoT and NextG networks [287, 288] and the Zero-Trust approach is one key direction the industry is moving towards. Zero-Trust architecture (ZTA) [289] is a framework that works on the underlying principle that no one accessing the network and its resources can be blindly trusted. The objective is to prevent data breaches and other cybersecurity attacks. While there are system checks in place such as authorizing users based on their unique ID and password, ZTA’s main function is to authorize each user’s requests on a case-by-case basis instead of authorizing the user. The key elements of ZTA are dynamic risk assessment and trust evaluation. Application of Artificial Intelligence has been significantly explored by researchers [290, 291] as well as intelligent ZTA for 6G networks [292, 293, 294]. Given the vast features of GNN, it promises to be a key enabler for ZTA-enabled systems. GNN can be used for Intelligent Network Security State Analysis (INSSA) [292], Trust Evaluation [295]. While there has been preliminary research done on GNN for Zero-Trust systems, it remains an exciting area of research towards 6G and NextG networks.
VIII Conclusion
In this article, we presented a comprehensive survey of GNN with a focus on the IoT and NextG networks. First, we provided a detailed description of GNN by starting with the common terminologies and their definitions before elaborating on its architecture. To complete this introduction to GNN, we provide detailed insight into the commonly used GNN variants. Second, we discussed the applications of GNN for IoT where we reviewed the existing literature focusing on IoT Network Intrusion Detection and IoT Data Fusion. Third, we looked into how GNNs have been leveraged for spectrum awareness and reviewed the existing literature focusing on RF signal classification and RF spectrum sensing. Fourth, we delved into applications of GNN for networking. Here, we discussed the existing literature focusing on applications of GNN for network and signal characteristics prediction, routing optimization, congestion control, mobile edge computing, digital twin networks, and unmanned aerial networks. Fifth, we focused on applications of GNN for tactical systems where we discussed existing literature on tactical communication and sensing systems, target recognition and localization. Finally, based on the detailed survey that we have conducted, we have listed the possible research directions for improving GNN features as well as open research areas that will benefit from GNN’s versatility.